OpenStudy (anonymous):

A spherical water balloon has a radius of 6 inches. How many cubic inches of water will it hold? a. 288ð in.3 b. 78ð in.3 c. 216ð in.3 d. 144ð in.3

hero (hero):

\[V_{\text{sphere}} = \frac{4\pi r^3}{3}\]

OpenStudy (anonymous):

@Hero I got 904.7808

OpenStudy (anonymous):

What does this symbol mean? ð

hero (hero):

pi I suppose.

OpenStudy (theeric):

I got the same volume formula and my volume on the windows calculator was \(904.7786842338604526772412943845\) using its estimation of \(pi\). The exact answer is \(288\ \pi\).

hero (hero):

288 pi = 904.7808

OpenStudy (theeric):

And so I agree with @Hero again. Maybe it was just an issue with the computer code. I don't see why anyone would substitute a symbol for \(\pi\) if that's what that means.

OpenStudy (theeric):

And the units are \([\text{in.}^2]\)

OpenStudy (theeric):

So, that looks right..

hero (hero):

cubic is in^3

OpenStudy (anonymous):

Thank you so much I get it now! =)

OpenStudy (theeric):

:)

OpenStudy (theeric):

Pfft, \([\text{in}^3]\), you're right again.